# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series

### v1

@@ 4.4 Seasonal and Trend Terms
Figure 1 and Supplementary figure S3 show varying levels of yearly and weekly seasonality in the currency data in different regions. There is a recurring yearly pattern driven by popular travel periods and a clear weekly seasonality reflecting variations between days of the week. Consequently, we include weekly and yearly seasonal variables and a trend variable in the predictors $\mathbf{X}_{t}$ and the same weekly and yearly seasonal variables in $\boldsymbol{z}_{t}$. Both $\phi_{t}$ and the elements of $\boldsymbol{\eta}_{t}$ are modeled with a linear trend and Fourier terms for seasonal variables, using pairs of $\left(\sin{\frac{2k\pi t}{w_{\text{season}}}},\cos{\frac{2k\pi t}{w_{\text{season}}}}\right)$ for $k=1,\dots,K_{\text{season}}\leq\frac{w_{\text{season}}}{2}$, with $w_{\text{season}}=7$ for weekly seasonality and $w_{\text{season}}=365.25$ for yearly seasonality. The orthogonality of the Fourier terms aids in model convergence. The B-DARCH, B-DARMA, and B-tVARMA all share the same $\mathbf{X}_{t}$ while the B-DARCH and B-DARMA share the same $\boldsymbol{z}_{t}$.
@@ 4.5 Model Specification and Validation
The models are fit using data from January 1, 2017, to June 30, 2020, as the training set. The validation set, spanning July 1, 2020, to September 30, 2020, is used to determine optimal lags and exogenous covariates. The test set from October 1, 2020, to December 31, 2020, is used for final model comparison.
For the initial specification, we fix $P=1$, $Q=0$, and $K_{\text{week}}=3$ for each region, allowing the intra-year Fourier terms to vary. The optimal number of Fourier terms ($K_{\text{year}}$) is selected based on the model’s performance on the validation set. With the optimal ($K_{\text{year}}$) established, we vary $P$ and $Q$ to identify the best-performing orders of the auto-regressive (AR) and moving average (MA) components on the validation set. Using these specifications, the models are retrained up until September 30, 2020, and tested on the independent test set for the B-DARMA, B-DARMA-DARCH, and B-tVARMA models.
4.6 Priors for B-DARCH, B-DARMA, and B-tVARMA
For all models (B-DARCH, B-DARMA, and B-tVARMA), we assume each component’s auto-regressive (AR) and moving average (MA) elements, $a_{rs}$ and $b_{rs}$, are strongest with their own lags. Thus, we assign weakly informative normal priors $N(.4,.5^{2})$ to $a_{rrp},b_{rrq}$ for diagonal elements ($r=1,\dots,4,q=1,\dots,3$), and $N(0,.5^{2})$ for non-diagonal elements in the $\boldsymbol{A}$ and $\boldsymbol{B}$ matrices.
The priors for the Fourier, trend, and intercept regression coefficients are $N(0,1)$, $N(0,.1^{2})$, and $N(0,2^{2})$ respectively. For the B-DARMA and B-DARCH, the precision model predictors $\boldsymbol{z}_{t}$ have the same seasonal terms as the mean models, with corresponding priors for the coefficients in $\boldsymbol{\gamma}$ matching those of the $\boldsymbol{\eta}$ terms. The B-DARCH has a $N(0,1)$ prior on its AR $\alpha$ and MA $\tau$ terms.
For the B-tVARMA, we impose a half normal prior $N^{+}(0,.5^{2})$ on the scale parameter $\sigma$ and the correlation matrix $\Omega$ has a $\text{LKJ}(3)$ prior on its Cholesky factor $L_{\Omega}$.
@@ 4.7.1 Model Specification
4.7.2 Performance on Test Set
The Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and major currency are shown in Table 2, while Figures 36 display the corresponding out‐of‐sample forecasts $\bar{\mu}_{tj}$. When errors are aggregated over all currencies within each region, the B-DARCH model consistently delivers the lowest total FMAE and FRSS. For example, in Region 1 the total FMAE (FRSS) for B-DARCH is 2.84 (3.15), which is markedly lower than 7.65 (19.11) for B-DARMA and 5.20 (10.62) for B-tVARMA. In Region 2, the advantage of B-DARCH is even more pronounced: its total FMAE is 13.55 (with an average of 2.26) and total FRSS is 73.93 (average 12.32), compared to 14.71 (2.45) and 81.87 (13.65) for B-DARMA, and 17.13 (2.86) and 104.43 (17.41) for B-tVARMA. Regions 3 and 4 exhibit similar aggregate advantages for B-DARCH.
At the individual-currency level, however, there are still several cases in which either B-DARMA or B-tVARMA yield lower FMAE—particularly in Regions 3 and 4. In Region 2, by contrast, the B-DARCH yields the smallest FMAE across all major currencies (CAD, CHF, EUR, GBP, USD, and “other”). In Regions 3 and 4, there remain a few currencies that favor one of the alternative models. For instance, in Region 3, B-DARMA produces slightly lower FMAE than B-DARCH for JPY (2.17 vs. 2.20) and NZD (2.91 vs. 3.57), while in Region 4, B-DARMA outperforms B-DARCH for EUR (0.58 vs. 1.29) but not for most of the other currencies. Notably, B-tVARMA is especially competitive for certain currencies in Region 1 (e.g., EUR), likely because its assumption of constant variance aligns well with periods of relatively stable volatility. Nonetheless, when the full basket of currencies is considered, the time-varying volatility framework of B-DARCH leads to the most accurate overall forecasts.
Supplementary Table S2 shows the empirical coverage rates of the 95% credible intervals for the forecasts across regions and currencies. The B-DARCH model exhibits relatively stable coverage that is generally closer to the nominal 95% level, with mean coverage rates of 0.92, 0.91, 0.87, and 0.91 in Regions 1, 2, 3, and 4, respectively. By contrast, the B-DARMA model shows greater variability, with mean coverages ranging from 0.82 in Region 1 down to 0.55 in Region 3, and 0.82 in Region 2 and 0.84 in Region 4 B-tVARMA tends to undercover, with mean rates of 0.63, 0.86, 0.69, and 0.85 across Regions 1–4. These results suggest that although B-tVARMA can occasionally yield competitive point forecasts, its assumption of time-invariant volatility often leads to overconfident interval estimates. In contrast, the dynamic volatility specification in B-DARCH not only improves point forecast accuracy but also delivers more reliable interval estimates.
@@ 4.7.2 Performance on Test Set
B-DARCH B-DARMA B-tVARMA B-DARCH B-DARMA B-tVARMA
1 CAD 0.55 0.87 1.21 0.49 1.07 1.75
1 EUR 0.61 2.19 0.55 0.53 5.23 0.55
1 GBP 0.18 0.43 0.21 0.05 0.22 0.06
1 JPY 0.09 0.18 0.06 0.01 0.04 0.01
1 USD 1.08 3.13 2.52 1.91 11.70 7.65
1 other 0.33 0.85 0.66 0.17 0.85 0.60
1 total 2.84 7.65 5.20 3.15 19.11 10.62
1 average 0.47 1.27 0.87 0.53 3.19 1.77
2 CAD 0.35 0.38 0.37 0.22 0.24 0.27
2 CHF 0.83 0.89 0.96 0.97 1.14 1.36
2 EUR 4.51 4.85 5.51 29.70 34.50 42.20
2 GBP 4.06 4.13 5.32 23.00 24.00 35.30
2 USD 1.28 1.68 1.68 3.04 4.59 4.60
2 other 2.52 2.78 3.29 17.00 17.40 20.70
2 total 13.55 14.71 17.13 73.93 81.87 104.43
2 average 2.26 2.45 2.86 12.32 13.65 17.41
3 AUD 6.29 8.24 7.78 57.50 88.50 80.80
3 EUR 0.63 0.93 0.87 0.52 1.09 0.99
3 JPY 2.20 2.17 2.34 23.70 24.10 22.70
3 NZD 3.57 2.91 3.57 14.60 9.96 14.70
3 USD 0.82 0.79 1.02 1.00 0.98 1.52
3 other 2.69 5.09 3.06 13.30 34.80 15.30
3 total 16.20 20.12 18.64 110.62 159.43 136.01
3 average 2.70 3.35 3.11 18.44 26.57 22.67
4 BRL 2.84 4.36 6.38 10.80 23.60 46.20
4 CLP 2.32 2.35 3.09 7.76 8.45 13.00
4 EUR 1.29 0.58 0.64 2.08 0.51 0.60
4 MXN 1.27 2.85 1.54 2.36 8.71 2.98
4 USD 2.70 4.14 2.94 8.68 19.60 9.93
4 other 0.79 0.97 1.94 0.87 1.46 4.93
4 total 11.21 15.25 16.53 32.55 62.33 77.64
4 average 1.87 2.54 2.76 5.43 10.39 12.94
@@
Figure 3: Airbnb data analysis - Region 1: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 4: Airbnb data analysis - Region 2: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 5: Airbnb data analysis - Region 3: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 6: Airbnb data analysis - Region 4: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
@@ 4.7.3 Residual Analysis
The partial autocorrelation function (PACF) of the sum of standardized squared residuals across all components on the test set was computed for each region and model (Supplementary Figures S9S12). All models exhibit a large peak at lag 1, reflecting immediate persistence in the residuals. However, the magnitude of this spike is noticeably smaller for the B-DARCH model. We consider a PACF value to be significant if it exceeds a threshold of 0.2; any autocorrelation values above this level indicate meaningful residual dependence.
Beyond lag 1, the B-DARCH model shows only one PACF value exceeding the 0.2 significance threshold, suggesting effective capture of short-term dependencies and minimal residual autocorrelation at higher lags. In contrast, both B-DARMA and B-tVARMA have PACF values surpassing the 0.2 threshold at multiple lags beyond lag 1, particularly between lags 2 and 10, indicating that residual dependencies persist across a broader range of lags in these models.
These suggest that the B-DARCH model achieves better residual independence and stability, with remaining residual structure potentially explained by unmodeled market-specific holidays, events, and policy implementations.
@@
A complete discussion of the posterior densities for the DARCH parameters (auto-regressive $\alpha$ and moving average $\tau$ coefficients) in each region appears in the Supplementary Material (Section Parameter Estimates). Briefly, Regions 2 and 3 exhibit higher $\alpha$ values indicating more persistent volatility, while Region 1 shows moderate persistence ($0.750.25<\alpha<$ 0.75), and Region 4 displays a more complex multi-lag structure. These region-specific patterns underscore the flexibility of our B-DARMA-DARCH approach to accommodate various volatility regimes.

### v2

@@ 4.4 Seasonal and Trend Terms
Figure 1 and Supplementary figure S3 show varying levels of yearly and weekly seasonality in the currency data in different regions. There is a recurring yearly pattern driven by popular travel periods and a clear weekly seasonality reflecting variations between days of the week. Consequently, we include weekly and yearly seasonal variables and a trend variable in the predictors $\mathbf{X}_{t}$ and the same weekly and yearly seasonal variables in $\mathbf{z}_{t}$. Both $\phi_{t}$ and the elements of $\boldsymbol{\eta}_{t}$ are modeled with a linear trend and Fourier terms for seasonal variables, using pairs of $\left(\sin{\frac{2k\pi t}{w_{\text{season}}}},\cos{\frac{2k\pi t}{w_{\text{season}}}}\right)$ for $k=1,\dots,K_{\text{season}}\leq\frac{w_{\text{season}}}{2}$, with $w_{\text{season}}=7$ for weekly seasonality and $w_{\text{season}}=365.25$ for yearly seasonality. The orthogonality of the Fourier terms aids in model convergence. The B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA all share the same $\mathbf{X}_{t}$ while the B-DARCH and B-DARMA share the same $\mathbf{z}_{t}$.
@@ 4.5 Model Specification and Validation
We train on January 1, 2017 to June 30, 2020; July 1 to September 30, 2020 serves as a validation window for choosing lags and seasonal complexity; October 1 to December 31, 2020 is the held-out test set. For the initial pass we set $P=1$, $Q=0$, and $K_{\text{week}}=3$ for each region, then select the optimal number of yearly Fourier terms $K_{\text{year}}$ by validation performance. With $K_{\text{year}}$ fixed, we vary $P$ and $Q$ on the validation set to choose AR/MA orders. Using the chosen specifications, all four models (B-DARMA, B-DARCH, B-tVARMA, and B-TVP-tVARMA) are refit through September 30, 2020 and evaluated on the independent test set.
4.6 Priors
Shared structure.
Across models we encode the belief that own–lag effects dominate by assigning diagonal AR/MA entries $(a_{rrp},b_{rrq})\sim\mathcal{N}(0.4,\,0.5^{2})$ and off–diagonals $\sim\mathcal{N}(0,\,0.5^{2})$. Seasonal Fourier terms, linear trends, and intercepts use $\mathcal{N}(0,1)$, $\mathcal{N}(0,0.1^{2})$, and $\mathcal{N}(0,2^{2})$, respectively. For Dirichlet models, the precision–side design $\mathbf{z}_{t}$ uses the same seasonal basis as the mean design $\mathbf{X}_{t}$.
B-DARMA and B-DARCH.
In both Dirichlet models the precision covariates $\mathbf{z}_{t}$ adopt the same priors as on the mean side. For B-DARCH, the AR and MA coefficients in $\log\phi_{t}$ each have $\mathcal{N}(0,1)$ priors, and the initial log‑precision is given a weakly informative normal prior.
B-tVARMA.
On the ALR scale we take
$\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}),\qquad\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0.5),\;L_{\boldsymbol{\Omega}}\sim\mathrm{LKJ\_Corr\_Cholesky}(3),$
yielding an unrestricted yet regularized covariance for the transformed-Gaussian likelihood.
B–TVP–tVARMA.
The TVP comparator uses the same ALR‑scale likelihood $\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\boldsymbol{\Sigma})$ but with $\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma})$, where each component $\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0,0.5^{2})$ and $L_{\Omega}\sim\mathrm{LKJ\_Corr\_Cholesky}(3)$. Time‑varying coefficient matrices follow mean‑reverting AR(1) evolutions around long‑run targets: $\mathbf{A}_{t,p}=\bar{\mathbf{A}}_{p}+\rho^{(A)}_{p}(\mathbf{A}_{t-1,p}-\bar{\mathbf{A}}_{p})+\tau^{(A)}_{p}\boldsymbol{Z}^{(A)}_{t,p}$ and analogously for $\mathbf{B}_{t,q}$, with $\operatorname{vec}(\boldsymbol{Z})\;\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\;\mathcal{N}(\mathbf{0},\mathbf{I})$, $\rho^{(A)}_{p},\rho^{(B)}_{q}\sim\operatorname{Beta}(9,1)$, and $\tau^{(A)}_{p},\tau^{(B)}_{q}\sim\mathrm{Half\!-\!Normal}(0,.1^{2})$. Long‑run targets are weakly centered by taking diagonal elements $\mathcal{N}(0.5,0.3)$ and off‑diagonals $\mathcal{N}(0,0.2)$, and initial states shrink to these targets with $\mathcal{N}(0,0.2)$ per entry. Exogenous effects are shared across models via $\boldsymbol{\beta}\sim\mathcal{N}(\mathbf{0},\beta_{\!sd}^{2}\mathbf{I})$.
@@ 4.7.1 Model Specification
Aggregate performance.
Table 4 reports Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and currency across four models: B-DARCH, B-DARMA, B-tVARMA, and a time-varying‑parameter VARMA fit in ALR space (TVP-tVARMA). The corresponding out-of-sample forecasts $\bar{\mu}_{tj}$ are shown in Figures 36.
When errors are aggregated over currencies within a region (the total rows of Table 4), B-DARCH attains the lowest total FMAE and FRSS in Regions 1–4. For Region 1, totals are FMAE/FRSS == 2.84/3.15 (B-DARCH) vs. 7.65/19.11 (B-DARMA), 5.20/10.62 (B-tVARMA), and 4.55/8.12 (B-TVP-tVARMA). For Region 2, totals are 13.55/73.93 (B-DARCH) vs. 14.71/81.87 (B-DARMA), 17.13/104.43 (B-tVARMA), and 14.68/90.49 (B-TVP-tVARMA). For Region 3 the ordering is 16.23/110.83 (B-DARCH) << 20.17/159.66 (B-DARMA) << 20.73/143.39 (B-TVP-tVARMA) << 23.70/187.09 (B-tVARMA). For Region 4, B-DARCH again leads with 11.21/32.55, followed by B-TVP-tVARMA (12.22/39.85), B-DARMA (15.25/62.33), and B-tVARMA (16.53/77.64). These results reinforce that explicitly modeling time‑varying volatility via DARCH yields the most accurate aggregate forecasts across regions. Across regions the B-TVP-tVARMA comparator required the longest runtime, followed by B-DARCH, while B-DARMA and B-tVARMA typically completed near one hour (Table S4).
Per-currency comparisons.
Currency-level minima vary by region. In Region 2, B-DARCH delivers the lowest FMAE for five of six currencies (CAD, CHF, EUR, USD, and other), with B‑TVP‑tVARMA best for GBP. In Region 3, B-DARCH is best for AUD, EUR, and other; B‑TVP‑tVARMA is best for JPY and NZD; and B-DARMA has the smallest USD FMAE. Region 4 shows a similar mix: B-DARCH leads for BRL, CLP, and other; B‑TVP‑tVARMA leads for MXN and USD; B-DARMA leads for EUR. In Region 1, B‑DARCH is best for CAD, USD, and other on both FMAE and FRSS; TVP‑tVARMA is best for EUR and JPY; and GBP is effectively a tie after rounding, with TVP‑tVARMA narrowly ahead on both metrics. Overall, the table indicates that the DARCH volatility component is beneficial across all regions, with especially large aggregate gains where volatility shifts are most pronounced.
Figure 3: Airbnb data analysis — Region 1: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 4: Airbnb data analysis — Region 2: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 5: Airbnb data analysis — Region 3: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 6: Airbnb data analysis — Region 4: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Table 3: Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) by region, currency, and model (values $\times 10^{2}$)
@@
DARCH DARMA tVARMA TVP-tVARMA DARCH DARMA tVARMA TVP-tVARMA
1 CAD 0.55 0.87 1.21 1.05 0.49 1.07 1.75 1.34
1 EUR 0.61 2.19 0.55 0.48 0.53 5.23 0.55 0.42
1 GBP 0.18 0.43 0.21 0.18 0.05 0.22 0.06 0.04
1 JPY 0.08 0.18 0.06 0.05 0.01 0.04 0.01 0.00
1 USD 1.08 3.13 2.52 2.21 1.91 11.70 7.65 5.85
1 other 0.33 0.85 0.66 0.57 0.17 0.85 0.60 0.46
1 total 2.84 7.65 5.20 4.55 3.15 19.11 10.62 8.12
1 average 0.47 1.27 0.87 0.76 0.53 3.19 1.77 1.35
2 CAD 0.35 0.38 0.37 0.46 0.22 0.24 0.27 0.36
2 CHF 0.83 0.89 0.96 1.19 0.97 1.14 1.36 2.00
2 EUR 4.51 4.85 5.51 5.65 29.70 34.50 42.20 49.70
2 GBP 4.06 4.13 5.32 2.83 23.00 24.00 35.30 14.90
2 USD 1.28 1.68 1.68 1.76 3.04 4.59 4.60 4.93
2 other 2.52 2.78 3.29 2.79 17.00 17.40 20.70 18.60
2 total 13.55 14.71 17.13 14.68 73.93 81.87 104.43 90.49
2 average 2.26 2.45 2.86 2.45 12.32 13.65 17.41 15.08
3 AUD 6.31 8.26 9.98 8.73 57.70 88.70 116.00 89.00
3 EUR 0.64 0.93 2.40 2.10 0.53 1.10 6.03 4.62
3 JPY 2.20 2.18 2.33 2.04 23.70 24.10 23.00 17.60
3 NZD 3.57 2.92 1.37 1.20 14.60 9.98 2.75 2.11
3 USD 0.82 0.79 2.48 2.17 1.00 0.98 6.61 5.06
3 other 2.69 5.09 5.14 4.49 13.30 34.80 32.70 25.00
3 total 16.23 20.17 23.70 20.73 110.83 159.66 187.09 143.39
3 average 2.71 3.36 3.95 3.46 18.47 26.61 31.18 23.90
4 BRL 2.84 4.36 6.38 3.42 10.80 23.60 46.20 15.80
4 CLP 2.32 2.35 3.09 2.46 7.76 8.45 13.00 9.78
4 EUR 1.29 0.58 0.64 1.88 2.08 0.51 0.60 4.15
4 MXN 1.27 2.85 1.54 1.15 2.36 8.71 2.98 1.98
4 USD 2.70 4.14 2.94 2.35 8.68 19.60 9.93 6.93
4 other 0.79 0.97 1.94 0.96 0.87 1.46 4.93 1.21
4 total 11.21 15.25 16.53 12.22 32.55 62.33 77.64 39.85
4 average 1.87 2.54 2.76 2.04 5.43 10.39 12.94 6.64
@@ 4.7.2 Residual Analysis
We compute the partial autocorrelation function (PACF) of the sum of standardized squared residuals (SSR) on the test set for each region and model (Supplementary Figures S9S12). All four models show a prominent spike at lag 1, but its magnitude is systematically smaller under B-DARCH. Using a practical significance threshold of $0.2$, B-DARCH exhibits only isolated exceedances beyond lag 1, indicating that short‑run volatility dynamics have largely been captured. By contrast, B-DARMA, B-tVARMA, and B-TVP-tVARMA display multiple exceedances between lags 2 and 10 in several regions, suggesting remaining residual dependence at intermediate lags. This pattern is consistent with the design of the DARCH component, which allows volatility to adapt dynamically to recent shocks and past variability, thereby reducing serial structure in SSR.
Supplementary Table S2 summarizes empirical coverage of 95% credible intervals. Mean coverage for B-DARCH remains close to nominal across regions (0.92, 0.91, 0.87, 0.91 in Regions 1–4). B-TVP-tVARMA’s coverage is mixed—strong in Region 4 (0.91) but under-covering in Region 1 (0.65) and Region 2 (0.75)—while B-DARMA varies widely and B-tVARMA often undercovers. Together with the forecast metrics, these results highlight that explicitly modeling time‑varying volatility improves both point accuracy and uncertainty calibration, especially in regions with pronounced volatility shifts.
@@
A complete discussion of the posterior densities for the DARCH parameters (auto-regressive $\alpha$ and moving average $\tau$ coefficients) in each region appears in the Supplementary Material (Section 6). Briefly, Regions 2 and 3 exhibit higher $\alpha$ values indicating more persistent volatility, while Region 1 shows moderate persistence ($0.25<\alpha<0.75$), and Region 4 displays a more complex multi-lag structure. These region-specific patterns underscore the flexibility of our B-DARCH approach to accommodate various volatility regimes.